OpenStudy (anonymous):

a chemist wants to obtain 80 liters of a solution that is 12% hyrdochloric acid.how many liters of 10% hydrochloric acid solution and 20% hyrochloric acid should be mixed? helpp going crazy

OpenStudy (kropot72):

Amount of hydrochloric acid in desired solution = 80 * 0.12 Let A = amount of 10% hydrochloric solution that is needed. Let B = amount of 20% hydrochloric solution that is needed. 0.1A + 0.2B = 80 * 0.12 = 9.6 liters ................(1) A + B = 80 liters...............................................(2) Now you just need to solve the system of equations (1) and (2) to find A and B.

OpenStudy (anonymous):

omg thank you so much!!

OpenStudy (kropot72):

You're welcome. Please let me know if you need help in solving the system of equations.

OpenStudy (anonymous):

im a bit confused

OpenStudy (anonymous):

what's the next step